# Jaime Angulo Pava  and  Alexander Muñoz

### v3

@@ 5.4. Extensions generating unitary dynamics parameterized by linear subspaces
Consider the space $Y_{0}=\mbox{span}\{(1,1,0)^{T}\}$. Note $\vec{\textbf{U}}\in Y_{0}$ if and only if $\phi(-L)=\phi(L)$. Also, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\phi^{\prime}(L)$. We therefore have that
$D(\mathcal{H}_{Y_{0}})=\{u\in D(\mathcal{H}_{0}^{*})\mid\ \phi(-L)=\phi(L)\ \mbox{and }\phi^{\prime}(L)=\phi^{\prime}(-L)\}.$
Example 10.
Consider the space $Y=\mbox{span}\{(1,0,-1)\}$. Consequently, $Y^{\perp}=\{(x,y,w)\mid x=w\}$. Note $\vec{\textbf{U}}\in Y$ if and only if $\phi(L)=0$ and $\phi(-L)=-\psi(L)$. Similarly, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\psi^{\prime}(L)$. According to Proposition 5.5, ${\mathcal{H}}_{Y}$ defines a self-adjoint extension of $\mathcal{H}_{0}$ with domain
$D(\mathcal{H}_{Y})=\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\phi(-L)=-\psi(L),\ \phi(L)=0,\ \mbox{and}\ \phi^{\prime}(-L)=\psi^{\prime}(L)\}.$ (5.25)
5.5. The Schrödinger operator on $\mathcal{T}$-shaped metric graphs
As highlighted in the introduction of the section, the analysis performed in the previous subsections for the Schrödinger equation on looping-edge graphs extends to $\mathcal{T}$-shaped graphs, i.e., to graphs $\mathcal{G}$ that consists on an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with 00 (see Figure 4). A similar frame-work can be employed to obtain self-adjoint extensions of $\mathcal{H}_{0}$ posed on these kind of metric graphs. We proceed to exhibit some examples of them.
$L$
00
Figure 4. $\mathcal{T}$-shaped metric graph.
Example 11.
Let $N=1$ (exactly one half-line attached to 00) and consider again the operator $\mathcal{L}$ defined in (5.18) with $m_{1}=1$ and $m_{2}=0$ (see (5.20)). Instead of $m_{3}=-1$, let us now consider $m_{3}=1$. It can be obtained analogously to (5.20) an extension of $\mathcal{H}_{0}$ with domain of functions in $H^{2}(\mathcal{G})$ such that
$\begin{split}&\psi^{\prime}(0+)=\phi^{\prime}(0-),\ \phi^{\prime}(-L)=0,\ \phi(0-)=\phi(-L)\ \\ &\mbox{ and }\phi(0-)-\psi(0+)=\frac{m_{4}}{2}\psi^{\prime}(0+),\end{split}$ (5.26)
where $0+$ and $0-$ refer to the right and left-handed limits respectively.
Thus, (5.26) is an example of the kind of domain of extensions that might be related to interactions of $\delta^{\prime}$-type in a $\mathcal{T}$-graph.
With the following example we intend to present, under this frame of reference, all the extensions generating unitary dynamics with continuity in the derivatives at the vertex. This family includes the so-called $\delta^{\prime}$ interactions. For simplicity we work with $N=1$ and $N=2$.
Example 12.
Let us first set $N=1$ and let $\mathcal{L}$ be defined by
$\mathcal{L}=\begin{pmatrix}2&m_{1}&0&m_{2}\\ 0&1&0&0\\ 0&m_{2}&2&m_{3}\\ 0&0&0&1\end{pmatrix}.$ (5.27)
It can be seen that in a respective boundary system $\mathcal{L}$ is a unitary operator which defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ via the relation $\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(0+)\end{pmatrix}=\begin{pmatrix}\partial\phi(0-)\\ \partial\phi(0-)\end{pmatrix}$. That is, an extension $\mathcal{H}_{\mathcal{L}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}(0+),\ \phi(0+)-2\phi(-L)=(m_{1}+m_{2})\psi^{\prime}(0+)\mbox{ and }\\ &\phi(0-)-2\psi(0+)=(m_{2}+m_{3})\psi^{\prime}(0+).\end{split}$ (5.28)
Adding the last two equations in (5.28) we get
$\phi(0-)-\phi(-L)=\psi(0+)+\frac{m_{1}+2m_{2}+m_{3}}{2}\psi^{\prime}(0+).$ (5.29)
In particular, if $m_{1}=-m_{3}$ it follows
$\phi(0-)-\phi(-L)=\psi(0+)+m_{2}\psi^{\prime}(0+).$ (5.30)
Similarly, let us assume $N=2$ and consider the operator
$\mathcal{L}_{2}=\begin{pmatrix}3&m_{1}&0&m_{2}&0&m_{3}\\ 0&1&0&0&0&0\\ 0&m_{2}&3&m_{4}&0&m_{5}\\ 0&0&0&1&0&0\\ 0&m_{3}&0&m_{5}&3&m_{6}\\ 0&0&0&0&0&1\end{pmatrix}$ (5.31)
It can be seen that $\mathcal{L}_{2}$ defines a unitary extension $H_{\mathcal{L}_{2}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions $(\phi,\psi_{1},\psi_{2})\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}_{1}(0+)=\psi^{\prime}_{2}(0+),\ \phi(0-)-3\phi(-L)=(m_{1}+m_{2}+m_{3})\psi^{\prime}(0+)\\ &\phi(0-)-3\psi_{1}(0+)=(m_{2}+m_{4}+m_{5})\psi^{\prime}(0+)\mbox{ and }\phi(0-)-3\psi_{2}(0+)=(m_{3}+m_{5}+m_{6})\psi^{\prime}(0+).\end{split}$ (5.32)
Adding the last three equations we get
$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+(m_{1}+m_{4}+m_{6}+2m_{2}+2m_{3}+2m_{5})\psi^{\prime}(0+).$ (5.33)
Of course there is a choice of the parameters that turn (5.33) into
$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.34)
Moreover, the operator $\mathcal{L}_{2}$ can be generalized to any amount of half-lines $N\in\mathbb{Z}^{+}$ by following a similar procedure to the one done in Example 7. That way it can be obtained an extension with domain made of functions satisfying continuity in the derivatives and
$\phi(0-)-\phi(-L)=\sum_{j=1}^{N}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.35)
@@ Acknowledgments
A. Muñoz is partially supported by CNPq, Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil grant 170246/2023-0 and INCTMat, Instituto Nacional de Ciência e Tecnologia de Matemática - Brazil. J. Angulo was partially funded by CNPq/Brazil Grant.
Data availability statement. The data that support the findings of this study are available upon request from the authors.

### v4

@@ 5.5. Extensions generating unitary dynamics parameterized by linear subspaces
Consider the space $Y_{0}=\mbox{span}\{(1,1,0)^{T}\}$. Note $\vec{\textbf{U}}\in Y_{0}$ if and only if $\phi(-L)=\phi(L)$ and $\psi(L)=0$. Also, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\phi^{\prime}(L)$. We therefore have that
[…]
5.6. The $\mathcal{T}$-shaped graph case
The abstract theory developed for the looping-edge graph $\mathcal{G}$ can be extended to address $\mathcal{T}$-shaped graphs (see Figure 2). These graphs are likewise characterized by a metric graph structure with $V=\{-L,L\}$ and $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$. In this setting, the vertex L-L is treated as an independent terminal node, and the looping condition $\phi(-L)=\phi(L)$ is not imposed.
Figure 2. A $\mathcal{T}$-shaped graph
Via the theory of boundary systems developed in Section 5.2, we characterize in the following example extension with a derivative coupling at the common vertex.
Example 10 (Extensions with prescribed continuity on the derivatives).
We seek all self-adjoint extensions of $\mathcal{H}_{0}$ whose domain includes the condition
$\phi^{\prime}(L)=\psi_{1}^{\prime}(L)=\cdots=\psi_{N}^{\prime}(L).$ (5.15)
From the boundary map definition and because of the sign in the second component, the conditions $\phi^{\prime}(L)=\psi_{j}^{\prime}(L)$ are equivalent to
$(\Gamma_{1}\mathbf{U})_{2}+(\Gamma_{1}\mathbf{U})_{j+2}=0,\qquad j=1,\ldots,N,$
or, setting $v_{j}:=e_{2}+e_{j+2}\in\mathcal{J}$,
$v_{j}^{\,*}\,\Gamma_{1}\mathbf{U}=0,\qquad j=1,\ldots,N.$ (5.16)
Since the prescribed conditions place no restriction on $\Gamma_{0}\mathbf{U}$, we take the full space $X=\mathcal{J}$ in Theorem 5.2. The domain condition then reads $\mathcal{L}\,\Gamma_{0}\mathbf{U}=\Gamma_{1}\mathbf{U}$ for all $\mathbf{U}\in D(\mathcal{H}_{X,\mathcal{L}})$. Substituting into (5.16),
$v_{j}^{\,*}\,\mathcal{L}\,\Gamma_{0}\mathbf{U}=0\qquad\text{for all }\Gamma_{0}\mathbf{U}\in\mathcal{J},$
which must hold structurally, i.e. as a condition on the rows of the matrix $\mathcal{L}\in M_{N+2}(\mathbb{C})$:
$\mathrm{row}_{j+2}(\mathcal{L})=-\,\mathrm{row}_{2}(\mathcal{L}),\qquad j=1,\ldots,N.$ (5.17)
We require $\mathcal{L}=\mathcal{L}^{*}$. The row constraints (5.17) propagate to column constraints via Hermiticity: $\mathrm{col}_{j+2}(\mathcal{L})=-\,\mathrm{col}_{2}(\mathcal{L})$ for $j=1,\ldots,N$. Enforcing both simultaneously and reading off the diagonal entries (which must be real), the most general Hermitian matrix satisfying (5.17) is
$\mathcal{L}=\begin{pmatrix}a&\beta&-\beta&\cdots&-\beta\\[4.0pt] \bar{\beta}&c&-c&\cdots&-c\\[2.0pt] -\bar{\beta}&-c&c&\cdots&c\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -\bar{\beta}&-c&c&\cdots&c\end{pmatrix},\qquad a,\,c\in\mathbb{R},\quad\beta\in\mathbb{C},$ (5.18)
where the lower-right $N\times N$ block is the constant matrix with every entry equal to $c$. This constitutes a four-real-parameter family $(a,c,\operatorname{Re}\beta,\operatorname{Im}\beta)$.
Expanding $\mathcal{L}\,\Gamma_{0}\mathbf{U}=\Gamma_{1}\mathbf{U}$ row by row and introducing the shorthand
$\sigma:=\phi(L)-\sum_{j=1}^{N}\psi_{j}(L),$
one obtains the following two independent conditions:
$\phi^{\prime}(-L)$ $=a\,\phi(-L)+\beta\,\sigma,$ (5.19)
$-\phi^{\prime}(L)$ $=\bar{\beta}\,\phi(-L)+c\,\sigma.$ (5.20)
The remaining $N$ equations (rows $3$ through $N+2$) each read $\psi_{j}^{\prime}(L)=\bar{\beta}\,\phi(-L)+c\,\sigma$, which is exactly -(5.20), thereby reproducing the prescribed condition (5.15) automatically.
Theorem 5.6.
The self-adjoint extensions of $\mathcal{H}_{0}$ whose domain satisfies (5.15) form the four-real-parameter family $\{\mathcal{H}_{a,c,\beta}\}_{a,c\in\mathbb{R},\,\beta\in\mathbb{C}}$, where
[…]
Remark 7.
Setting $\beta=0$ decouples the left endpoint from the vertex: (5.19) becomes a standard Robin condition $\phi^{\prime}(-L)=a\,\phi(-L)$, and (5.20) becomes a Robin condition at the vertex controlled by $c$ and the combination $\sigma$. More precisely, for parameters $Z_{1}\in\mathbb{R}$, $0\neq Z_{2}\in\mathbb{R}$ we have the self-adjoint extension $(H_{Z_{1},Z_{2}},D(H_{Z_{1},Z_{2}}))$ with
[…]
The choice $a=c=\beta=0$ yields the pure Neumann extension $\phi^{\prime}(-L)=0$, $\phi^{\prime}(L)=\psi_{j}^{\prime}(L)=0$ for all $j$.
5.7. A simple application to instability of standing waves of the cubic NLS
Let $D_{\mathrm{ext}}=D(\mathcal{H})$ where $D(\mathcal{H})$ is defined in (5.9) with $M^{\prime}=0$. That is, $D_{\mathrm{ext}}$ defines a self-adjoint extension of the Laplacian inducing pure periodic conditions on the loop and independent pure Neumann conditions on each half-line.
We seek real-valued standing waves $\mathbf{U}(t,x)=e^{i\omega t}\boldsymbol{\Phi}_{\omega}(x)$ with $\boldsymbol{\Phi}_{\omega}=(\phi_{\omega},\psi_{1,\omega},\dots,\psi_{N,\omega})\in D_{\mathrm{ext}}$. Substituting into (1.1) with $p=1$, each component of $\boldsymbol{\Phi}_{\omega}$ satisfies an independent stationary NLS equation $f^{\prime\prime}+f^{3}-\omega f=0$ on its respective edge.
For the loop component we consider a positive, real, $2L$-periodic solution constructed via the Jacobi dnoidal function. By Theorem 2.1 of [9], for $\omega>\omega^{*}:=\tfrac{\pi^{2}}{2L^{2}}$, the function
$\phi_{\omega}(x)=\eta_{1}\,\mathrm{dn}\!\left(\frac{\eta_{1}}{\sqrt{2}}\,x;\,k\right)$ (5.21)
with elliptic modulus $k\in(0,1)$ determined by
$k^{2}(\eta,\omega)=\frac{2\omega-2\eta^{2}_{2}}{2\omega-\eta^{2}_{2}},\qquad\eta^{2}_{1}+\eta^{2}_{2}=2\omega,\qquad 0<\eta_{2}<\eta_{1},$ (5.22)
solves the stationary loop equation. The half-line stationary equation is solved by the half-soliton
$\psi_{\omega}(x)=\sqrt{2\omega}\,\operatorname{sech}\!\bigl(\sqrt{\omega}(x-L)\bigr),\qquad x\in[L,+\infty),$
which is the restriction to $L,+\infty)$ of the full soliton centred at $x=L$. The Neumann condition $\psi_{\omega}^{\prime}(L)=0$ holds automatically because the full soliton is even about $L$. For $j=1,\dots,n\leqslant N$ we therefore set $\psi_{j,\omega}\equiv\psi_{\omega}$, and for $j=n+1,\dots,N$ we set $\psi_{j,\omega}\equiv 0$. The standing wave profile under study is thus (see Figure 3)
$\boldsymbol{\Phi}_{\omega,n}\equiv\boldsymbol{\Phi}_{\omega}=\bigl(\phi_{\omega},\,\underbrace{\psi_{1,\omega},\dots,\psi_{n,\omega}}_{n\text{ half-solitons}},\,\underbrace{0,\dots,0}_{N-n\text{ zero-solitons}}\bigr)\in D_{\mathrm{ext}},\qquad\omega>\omega^{*}.$ (5.23)
Figure 3. Schematic of the standing wave profile $\Phi_{\omega,n}$ with $N=3$ and $n=2$.
Theorem 5.7.
Let $L>0$ and $N\geqslant 1$. For every $n\in\{1,\dots,N\}$ and every $\omega>\omega_{*}$, the orbit
$\mathcal{O}(\mathbf{\Phi}_{\omega})=\bigl\{e^{i\theta}\bigl(\phi_{\omega}(\cdot+s),\,\psi_{1,\omega},\dots,\psi_{n,\omega},0,\dots,0\bigr):\theta\in\mathbb{R},\;s\in\mathbb{R}/(2L\mathbb{Z})\bigr\}$
is orbitally unstable in $H^{1}(\mathcal{G})$.
Proof.
The decoupled structure of $D_{\mathrm{ext}}$ reduces global well-posedness on $H^{1}(\mathcal{G})$ to independent problems on each edge. On the loop [L,L][-L,L] with periodic boundary conditions, global well-posedness in $H^{1}$ is classical and follows from semigroup theory together with conservation of energy and charge. On each half-line $L,+\infty)$ with Neumann boundary condition $\psi^{\prime}(L)=0$, well-posedness is reduced to the full-line case via the reflection method: extending $\psi$ to an even function on $\mathbb{R}$ embeds the problem into the cubic NLS on $\mathbb{R}$ restricted to even initial data, which is invariant under the flow. Global solutions for all $t\in\mathbb{R}$ are then obtained from the conservation of energy and charge via the Gagliardo–Nirenberg inequality.
Fix $\varepsilon>0$ and define the initial datum
$\mathbf{U}_{\varepsilon}(0):=\bigl(\phi_{\omega},\;\psi_{1,\omega+\varepsilon},\;\psi_{2,\omega},\dots,\psi_{n,\omega},\;0,\dots,0\bigr),$
where $\psi_{1,\omega+\varepsilon}(x)=\sqrt{2(\omega+\varepsilon)}\,\operatorname{sech}\!\bigl(\sqrt{\omega+\varepsilon}(x-L)\bigr)$ is the half-soliton at the shifted frequency $\omega+\varepsilon$. Since $\omega\mapsto\psi_{1,\omega}$ is $C^{1}$ as a map into $H^{1}([L,+\infty))$, the mean value inequality in Banach spaces gives
$\|\mathbf{U}_{\varepsilon}(0)-\mathbf{\Phi}_{\omega}\|_{H^{1}(\mathcal{G})}=\|\psi_{1,\omega+\varepsilon}-\psi_{1,\omega}\|_{H^{1}([L,+\infty))}\leqslant C_{\omega}\,\varepsilon,$ (5.24)
where $C_{\omega}=\sup_{\omega^{\prime}\in(\omega,\,\omega+1)}\|\partial_{\omega}\psi_{1,\omega^{\prime}}\|_{H^{1}([L,+\infty))}<\infty$. In particular, $\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(0),\mathcal{O}(\mathbf{\Phi}_{\omega}))\to 0$ as $\varepsilon\to 0$.
Since every component of $\mathbf{U}_{\varepsilon}(0)$ is an exact stationary profile at its respective frequency, each component satisfies its own NLS equation with its boundary conditions. By the uniqueness of $H^{1}$ solutions, the global solution is therefore given by
$\mathbf{U}_{\varepsilon}(t)=\bigl(e^{i\omega t}\phi_{\omega},\;e^{i(\omega+\varepsilon)t}\psi_{1,\omega+\varepsilon},\;e^{i\omega t}\psi_{2,\omega},\;\dots,\;e^{i\omega t}\psi_{n,\omega},\;0,\dots,0\bigr).$ (5.25)
At the time $t_{\varepsilon}:=\pi/\varepsilon$ one has $e^{i(\omega+\varepsilon)t_{\varepsilon}}=e^{i\omega t_{\varepsilon}}\cdot e^{i\pi}=-e^{i\omega t_{\varepsilon}}$, so the solution becomes
$\mathbf{U}_{\varepsilon}(t_{\varepsilon})=e^{i\omega t_{\varepsilon}}\bigl(\phi_{\omega},\;-\psi_{1,\omega+\varepsilon},\;\psi_{2,\omega},\dots,\psi_{n,\omega},\;0,\dots\bigr),$ (5.26)
that is, the first half-line component has flipped sign relative to all others.
We now derive a lower bound on the distance from the orbit at time $t_{\varepsilon}$. Take any orbit element $\mathbf{V}=e^{i\theta}(\phi_{\omega}(\cdot+s),\psi_{1,\omega},\dots)\in\mathcal{O}(\mathbf{\Phi}_{\omega})$ and set $\mu:=\theta-\omega t_{\varepsilon}$. Since the components are supported on disjoint edges and $\|(1-e^{i\mu})\psi_{j,\omega}\|_{L^{2}}^{2}\geqslant 0$ for $j=2,\dots,N$, we obtain
$\begin{split}\|\mathbf{U}_{\varepsilon}(t_{\varepsilon})-\mathbf{V}\|_{L^{2}(\mathcal{G})}^{2}&\geqslant\|\phi_{\omega}-e^{i\mu}\phi_{\omega}(\cdot+s)\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}+e^{i\mu}\psi_{1,\omega}\|_{L^{2}}^{2}\\ &=A+2\cos(\mu)\bigl(P-q(s)\bigr),\end{split}$ (5.27)
where
$\begin{split}A&:=2\|\phi_{\omega}\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2},\\ P&:=\langle\psi_{1,\omega+\varepsilon},\psi_{1,\omega}\rangle_{L^{2}([L,+\infty))},\\ q(s)&:=\langle\phi_{\omega},\phi_{\omega}(\cdot+s)\rangle_{L^{2}([-L,L])}.\end{split}$
Since $\cos\mu\in[-1,1]$, minimising the right-hand side of (5.27) over $\mu$ and then taking the infimum over $(\theta,s)$ gives
$\inf_{(\theta,s)}\|\mathbf{U}_{\varepsilon}(t_{\varepsilon})-\mathbf{V}\|_{L^{2}}^{2}\geqslant A-2\sup_{s\in\mathbb{R}/(2L\mathbb{Z})}|P-q(s)|.$ (5.28)
It is shown in Appendix A that
$A-2\sup_{s}|P-q(s)|\geqslant\min\!\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)=:c_{0}^{2}>0,$
where $q_{\min}:=\min_{s}q(s)>0$ and $c_{0}$ depends only on $\omega$, $L$, and $k$, but not on $\varepsilon$. Combining with (5.28) we conclude
$\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(t_{\varepsilon}),\,\mathcal{O}(\mathbf{\Phi}_{\omega}))\geqslant\mathrm{dist}_{L^{2}}(\mathbf{U}_{\varepsilon}(t_{\varepsilon}),\,\mathcal{O}(\mathbf{\Phi}_{\omega}))\geqslant c_{0}>0.$ (5.29)
Finally, given $\delta>0$ arbitrary, one can choose $\varepsilon>0$ small enough that $C_{\omega}\varepsilon<\delta$. Then (5.24) gives $\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(0),\mathcal{O}(\mathbf{\Phi}_{\omega}))<\delta$, while (5.29) gives $\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(t_{\varepsilon}),\mathcal{O}(\mathbf{\Phi}_{\omega}))\geqslant c_{0}$. This is precisely the opposite of orbital stability, and the proof is complete. ∎
@@ Acknowledgments
The authors express their gratitude to Prof. Nataliia Goloshchapova (IME-USP) for her invaluable comments on boundary systems, which significantly improved the results presented in the second part of this manuscript.
This study is financed, in part, by the São Paulo Research Foundation (FAPESP), Brazil. Process number 2024/20623-7.
A. Muñoz was partially supported by CNPq, Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil grant 170246/2023-0 and INCTMat, Instituto Nacional de Ciência e Tecnologia de Matemática - Brazil. J. Angulo was partially funded by CNPq/Brazil Grant.
Appendix A Lower bound computation
In this appendix we establish the inequality
$A-2\sup_{s\in\mathbb{R}/2L\mathbb{Z}}|P-q(s)|\geqslant\min\!\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)=:c_{0}^{2}>0,$
used in Section 5.7.
Since $\phi_{\omega}$ and $\psi_{\omega}$ are strictly positive, both $P$ and $q(s)$ are strictly positive. The map $s\mapsto q(s)$ is continuous on the compact set $\mathbb{R}/2L\mathbb{Z}$, so it attains strictly positive extrema $0<q_{\min}<q_{\max}$. The maximum is attained at $s=0$, with $q_{\max}=q(0)=\|\phi_{\omega}\|^{2}_{L^{2}}$, as a consequence of the Cauchy–Schwarz inequality and the invariance of the $L^{2}$ norm under periodic translation. Since $P$ is independent of $s$, the convexity of the absolute value gives
$\sup_{s}|P-q(s)|\leqslant\max\bigl(|P-q_{\min}|,\;|P-q_{\max}|\bigr).$
We argue in three cases according to the position of $P$ relative to $q_{\min},q_{\max}]$.
Case 1: $P\leqslant q_{\min}$. Since $q(s)\geqslant q_{\min}\geqslant P$ for all $s$, one has |Pq(s)|=q(s)P|P-q(s)|=q(s)-P and the supremum is attained at $s=0$:
$\sup_{s}|P-q(s)|=q_{\max}-P=\|\phi_{\omega}\|^{2}_{L^{2}}-P.$
Substituting back and cancelling the loop norm terms:
$\begin{split}A-2\sup_{s}|P-q(s)|&=2\|\phi_{\omega}\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2\|\phi_{\omega}\|_{L^{2}}^{2}+2P\\ &=\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}+2P.\end{split}$
All three terms are non-negative and $\|\psi_{1,\omega}\|_{L^{2}}^{2}=2\sqrt{\omega}$, so
$A-2\sup_{s}|P-q(s)|\geqslant 2\|\psi_{1,\omega}\|_{L^{2}}^{2}=4\sqrt{\omega}>0.$
Case 2: $q_{\min}\leqslant P\leqslant q_{\max}$. Here $|P-q_{\min}|=P-q_{\min}$ and $|P-q_{\max}|=q_{\max}-P$, so
$\sup_{s}|P-q(s)|=\max(P-q_{\min},\;q_{\max}-P).$
Subcase 2a: $q_{\max}-P\geqslant P-q_{\min}$. The supremum equals $q_{\max}-P=\|\phi_{\omega}\|_{L^{2}}^{2}-P$, and the same computation as in Case 1 yields
$A-2\sup_{s}|P-q(s)|=\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}+2P\geqslant 4\sqrt{\omega}>0.$
Subcase 2b: $P-q_{\min}>q_{\max}-P$. The supremum equals $P-q_{\min}$. Since $P\leqslant q_{\max}=\|\phi_{\omega}\|_{L^{2}}^{2}$, we have $2\|\phi_{\omega}\|_{L^{2}}^{2}-2P\geqslant 0$, and therefore
$\begin{split}A-2\sup_{s}|P-q(s)|&=2\|\phi_{\omega}\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2P+2q_{\min}\\ &\geqslant\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}+2q_{\min}\geqslant 2q_{\min}>0.\end{split}$
Combining both subcases, in Case 2:
$A-2\sup_{s}|P-q(s)|\geqslant\min\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)>0.$
Case 3: $P>q_{\max}$. Since $P>q(s)$ for all $s$, one has |Pq(s)|=Pq(s)|P-q(s)|=P-q(s) and the supremum is attained at the minimizer of $q$:
$\sup_{s}|P-q(s)|=P-q_{\min}.$
By the Cauchy–Schwarz inequality, $P\leqslant\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}\|\psi_{1,\omega}\|_{L^{2}}$, hence
$\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2P\geqslant\bigl(\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}-\|\psi_{1,\omega}\|_{L^{2}}\bigr)^{2}\geqslant 0.$
Therefore:
$\begin{split}A-2\sup_{s}|P-q(s)|&=2\|\phi_{\omega}\|_{L^{2}}^{2}+2q_{\min}+\bigl(\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2P\bigr)\\ &\geqslant 2\|\phi_{\omega}\|_{L^{2}}^{2}+2q_{\min}\geqslant 2q_{\min}>0.\end{split}$
Collecting all cases, we conclude that
$A-2\sup_{s\in\mathbb{R}/2L\mathbb{Z}}|P-q(s)|\geqslant\min\!\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)=:c_{0}^{2}>0,$
where $c_{0}^{2}$ depends only on $\omega$, $L$, and the elliptic modulus $k=k(\omega)$, but not on $\varepsilon$.